# Maintaining Adversarial Robustness in Continuous Learning

### v1

@@ 3.2 Double Gradient Projection (DGP)
The fundamental principle of our algorithm is concise: stabilizing the smoothed sample gradients (some implementation details are elaborated in the preceding subsection). The overall algorithmic flow is outlined as follows: Firstly, the neural network is trained on task $\mathcal{T}_{t}$ using the loss function that includes a penalty term of sample gradients, as shown in Eq. 1. The weight update is projected to be orthogonal to the base vectors in pool $\mathcal{P}$ if the sequential number $t>1$. Subsequently, after training, SVD is performed on the layer-wise inputs $\mathbf{X}^{l}_{t}$ to obtain base vectors that span the first gradient subspaces crucial for stabilizing the final outputs of the neural network on task $\mathcal{T}_{t}$. Lastly, another SVD is performed on the gradients of the layer-wise inputs with respect to the samples $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ to obtain base vectors that span the second gradient subspaces crucial for stabilizing the gradients of final outputs with respect to the samples on task $\mathcal{T}_{t}$. Note that in order to eliminate the redundancy between new bases and existing bases in the pool $\mathcal{P}$, both $\mathbf{X}^{l}_{t}$ and $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ are projected orthogonally onto $\mathcal{P}^{l}$ prior to performing SVD. Please refer to [4] for details. A compact pseudo-code of our algorithm is presented in Alg. 1.
@@
Input: Training dataset $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$ for task $\mathcal{T}_{t}\in\left\{\mathcal{T}_{1},\mathcal{T}_{2},\dots\right\}$, regularization strength $\lambda$, learning rate $\alpha$
@@
3:        Sample a batch from $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$
4:        Calculate $\left\{\nabla_{\mathbf{W}^{l}}\mathcal{L}\right\}$ with the loss $\mathcal{L}$ in Eq. 1
5:        if $t>1$ then
6:           $\nabla_{W^{l}}\mathcal{L}=\nabla_{W^{l}}\mathcal{L}-\left(\nabla_{W^{l}}\mathcal{L}\right)\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}$ $\forall l=1,2,...,L$triangleright Gradient projection for each layer
7:        end if
8:        $\mathbf{W}^{l}\leftarrow\mathbf{W}^{l}-\alpha\nabla_{\mathbf{W}^{l}}\mathcal{L}$triangleright Weight update
9:     end while
10:     $\mathbf{X}^{l}_{t}\leftarrow\mathbf{X}^{l}_{t}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\mathbf{X}^{l}_{t}$triangleright Ensure uniqueness of new bases
11:     Perform SVD on $\mathbf{X}^{l}_{t}$ and put bases into $\mathcal{P}^{l}$triangleright Construct first gradient subspace
12:     $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}\leftarrow\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}$
13:     Perform SVD on $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ and put bases into $\mathcal{P}^{l}$triangleright Construct second gradient subspace
14:  end for
@@ 4.1 Step
Baselines. For continual learning [9], we adopt (1) SGD, which serves as a naive baseline using stochastic gradient descent to optimize the neural network, (2) EWC [10], (3) SI [11], (4) GEM [12], (5) A-GEM [13], (6) OGD [1] and (7) GPM. For adversarial robustness, we adopt (8) IGR, and (9) Distillation [14].
We combine the methods from fields of continual learning (1$\sim$7) and adversarial robustness (8$\sim$9), such as EWC + IGR, to establish the baselines for continual robust learning. Additionally, we apply the FGSM [15], PGD [16], and AutoAttack [17] to generate adversarial attacks.
Figure 3: ACC varying with the number of learned tasks on datasets of Permuted MNIST (first row), Rotated MNIST (second row), CIFAR100 (third row) and miniImageNet (fourth row). ACC is measured on adversarial samples generated by AutoAttack (first column), PGD (second column) and FGSM (third column), as well as original samples (fourth column). The horizontal axis indicates the number of tasks learned by the neural network at present. The defense strategy used here is IGR. Errors bars denote standard deviation.
@@
Benchmarks. We evaluate our approach on four supervised benchmarks. Permuted MNIST and Rotated MNIST are variants of MNIST dataset of handwritten digits with 10 tasks applying random permutations of the input pixels and random rotations of the original images respectively [18, 19]. Split-CIFAR100 [11] is a random division of CIFAR100 into 10 tasks, each with 10 different classes. Split miniImageNet is a random division of a subset of the original ImageNet dataset [20] into 10 tasks, each with 5 different classes.
@@
The values of $\frac{\partial\mathbf{x}^{2}}{\partial\mathbf{x}}$ (initial term in Eq.16) are solely determined by the weights of the first layer when feeding the same samples (see Fig.1). There are two options to make $\frac{\partial\mathbf{x}^{2}_{p,t}}{\partial\mathbf{x}_{p}}=\frac{\partial\mathbf{x}^{2}_{p,p}}{\partial\mathbf{x}_{p}}$: fixing the first layer after learning task $p$ or assigning an independent first layer to each task. The latter option is chosen for all experiments, as the former seriously diminishes the neural network’s learning ability in subsequent tasks. To ensure fairness, the same setup is applied to the baselines. Further details on architectures can be found in Appendix B.2.
Training details: For the MNIST experiments, the batch size, number of epochs, and input gradient regularization $\lambda$ are set to 32/10/50, respectively. In the Split-CIFAR100 experiments, the values are 10/100/1, and for the Split-miniImageNet experiments, they are 10/20/1. SGD is used as the optimizer. All reported results are averaged over 5 runs with different seeds. The hyperparameter configurations for algorithms of adversarial attack and continuous learning are provided in Appendix B.3.
@@ 4.2.1 Adversarial Robustness
The robustness results on various datasets are presented in left three columns of Fig. 3. Under attacks with different strengths (AutoAttack $>$ PGD $>$ FGSM), the proposed approach (orange lines) consistently exhibits a high level of effectiveness in preserving the robustness of neural networks during continuous learning. In contrast, the baseline such as IGR+GEM (purple lines), which performs well on MNIST datasets against PGD and FGSM attacks, demonstrates a significant decrease when fronted with AutoAttack. The advantage of our approach becomes even more evident when the number of learned tasks increases.
In Fig. 3, the different color lines represent baselines combined by IGR with various methods of continuous learning. Except for the specifically tailored DGP, these methods do not maintain robustness well gained by IGR when the neural network parameters change to learn new tasks. To provide a comprehensive comparison, we include additional experimental results in Fig. 4, focusing on combinations of Distillation with methods of continuous learning. Distillation is another well-known defense method [21] that involves training two models - a teacher model is trained using one-hot ground truth labels and a student model is trained using the softmax probability outputs of the teacher model. The result demonstrates the model’s robustness gained by Distillation is also not effectively maintained against AutoAttack and PGD attacks after learning a sequence of new tasks.
@@
Considering the collective insights presented in Figs. 3 and 4, it is crucial to underscore that the pursuit of an effective defense demands a specially tailored algorithm adept at accommodating variations in model parameters. Mere combinations of existing defense strategies and continual learning methods, as demonstrated in our experiments, fall short of achieving the desired goal of continuous robustness. Our work contributes to the exploration of this under-researched and challenging research direction.
Moreover, there is a notable trend in Fig.4d: the blue line (representing the performance of Distill+GPM on original samples) exhibits a more rapid decline compared to the corresponding blue line in the fourth subplot of the first row in Fig.3 (representing IGR+GPM). Additionally, the purple and blue lines in Fig.4d (representing Distill+GEM and Distill+GPM) closely align with the green line in Fig.4d (representing Distill+SGD). These observations suggest incorporating Distillation into the training procedure compromise the efficacy of these continual learning methods.
Figure 4: As Fig. 3, but for the defense strategy Distillation on PMNIST dataset. Continual learning methods GEM and GPM are selected to combine with Distillation here, which show superior ACC in Fig. 3 compared to other baselines.
4.2.2 Performance
@@
We also assess the ability of the proposed approach for continual learning (ACC on original samples), as illustrated in the fourth column of Fig. 3. Our DGP algorithm demonstrates comparable performance to GPM and GEM on datasets of Permuted MNIST and Rotated MNIST, effectively addressing the issue of catastrophic forgetting on these two datasets. However, results on the datasets of Split-CIFAR indicate that the performance of DGP is slightly inferior to GDP. We speculate that the reason for this could be that DGP stores a larger number of bases after each task than GPM, as DGP constrains the weight updates to be orthogonal to two crucial subspaces – one for stabilizing the final output (required in GPM) and another for stabilizing the sample gradients. Orthogonality to more base vectors restricts weight updates to a narrower subspace, thereby limiting the plasticity of the model. Overall, our approach effectively maintains adversarial robustness while exhibiting continual learning ability.
In addition, it is noteworthy that the performance curves of several classical continual learning methods closely approximate that of naive SGD (green lines). This observation suggest again potential incompatibility between defense strategies (here IGR, i.e., sample gradient penalty) and continuous learning methods. Additional experimental results are provided in Appendix B.1.
@@ 4.2.3 Stabilization of sample gradients
Our approach maintains adversarial robustness by stabilizing the smoothed sample gradients. To valid its stabilization effect, we record the variation of sample gradients on the first task during continuous learning process. Specifically, we randomly select $n$ samples at the end of learning $\mathcal{T}_{1}$ and compute their gradients related to correspondingly final outputs. After learning each new task $\mathcal{T}_{t}$, we recompute their gradients. The variation of gradients between $\mathcal{T}_{1}$ and $\mathcal{T}_{t}$ is quantified by similarity measure:
@@
where $\mathbf{g}$ is a flattened vector of sample gradients. The results on various datasets are presented in Fig. 5. The orange line (representing DGP) shows a relatively flat downward trend, demonstrating the proposed approach indeed has the effect of stabilizing the sample gradients of previous tasks, even after weights of the neural network have been updated.
5 Related Work
Figure 5: Gradient variation of samples from the first task $\mathcal{T}_{1}$ during continuous learning process. The variations are quantified through similarity.
Continual Learning: Optimization. This class of continual learning approaches design and manipulate the optimization programs explicitly, often employing gradient projection as a typical strategy [22]. Replay-based methods like GEM [12] and A-GEM [13] impose constraints on weight updates to align with the direction of experience replay, thereby preserving previous input space and gradient space through historical training samples. Without saving historical training samples, OWM [23] and AOP [24] modify weight updates to the orthogonal direction of the input space of previous tasks. OGD [1] stores the gradient directions of previous tasks and orthogonalizes the current gradient (the gradients of all parameters are combined into a single vector) with respect to them. FS-DGPM [25] dynamically releases unimportant bases of GPM to improve learning plasticity and encourages the convergence towards a flat loss landscape. In contrast to OGD and GPM, Adam-NSCL [5] directly projects candidate weight updates into the unimportant gradient subspace, i.e., the current null space of the uncentered feature covariance from the previous tasks.
Adversarial Robustness: Regularization. As several attack algorithms rely on the model’s gradient to estimate local optima perturbations that can deceive the classifier, a stream of research has focused on adversarial robust optimization by applying regularization approaches to mitigate the impact of small input perturbations on output decisions, such as the spectral norm of the model’s Jacobian matrix [26], the spectral norm of the weight matrix [27] and the generated adversarial perturbation [28].
Our research does not focus on innovation in the single field of continual learning or adversarial robustness, but rather explores a novel direction that requires their integration. It expands the investigation of adversarial robustness to continuous learning scenarios, which holds significant value for enhancing the security and credibility of artificial intelligence systems in their adaptive evolution across multiple tasks or data domains.
@@ 6 Limitation and Discussion
In this work, we observe that the adversarial robustness gained by sophisticated defense strategies is easily erased when the neural network learns new knowledge. The direct combinations of existing defense strategies and continuous learning methods fail to effectively address this issue. In the context of continuous learning, a tailored algorithm is necessary to maintain the robustness brought originally by defense strategies. Therefore, we propose Double Gradient Projection, an approach that can prevents the adversarial robustness from rapidly degradation in the face of changes in model parameters.
From the experimental results, our approach has a limitation: as the number of base vectors becomes large, the stability of the neural network is enhanced to hold both robustness and performance across previous tasks. However, the plasticity of the neural network is restricted, which potentially reduces its learning ability for new tasks. Besides, the memory and computational time of our approach are double that of GPM, due to performing SVD twice after each learned task. If there are numerous tasks and the matrix consisting of orthogonal bases reaches full rank, we could approximate this matrix by performing SVD and selecting column vectors corresponding to a fraction of the largest singular values as the new orthogonal bases to free up the rank space.

### v2

@@ 3.2 Double Gradient Projection (DGP)
The fundamental principle of our algorithm is concise: stabilizing the smoothed sample gradients (some implementation details are elaborated in the preceding subsection). The overall algorithmic flow is outlined as follows: Firstly, the neural network is trained on task $\mathcal{T}_{t}$ with a class of defense algorithms through sample gradient smoothing. The weight update is projected to be orthogonal to all the base vectors in pool $\mathcal{P}$ if the sequential number $t>1$. Subsequently, after training, SVD is performed on the layer-wise outputs $\mathbf{X}^{l}_{t}$ to obtain base vectors for stabilizing the final outputs of the neural network on task $\mathcal{T}_{t}$. Lastly, another SVD is performed on the gradients of the layer-wise outputs with respect to the samples $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ to obtain base vectors for stabilizing the gradients of final outputs with respect to the samples on task $\mathcal{T}_{t}$. Note that in order to eliminate the redundancy between new bases and existing bases in the pool $\mathcal{P}$, both $\mathbf{X}^{l}_{t}$ and $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ are projected orthogonally onto $\mathcal{P}^{l}$ prior to performing SVD. A compact pseudo-code of our algorithm is presented in Alg. 1.
@@
Input: Training dataset $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$ for task $\mathcal{T}_{t}\in\left\{\mathcal{T}_{1},\mathcal{T}_{2},\dots\right\}$, regularization strength $\lambda$ and learning rate $\alpha$
@@
3:        Sample a batch from $\left\{\mathbf{X}_{t},\mathbf{Y}_{t}\right\}$ and Calculate $\left\{\nabla_{\mathbf{W}^{l}}\mathcal{L}\right\}$
4:        if $t>1$ then
5:           $\nabla_{W^{l}}\mathcal{L}=\nabla_{W^{l}}\mathcal{L}-\left(\nabla_{W^{l}}\mathcal{L}\right)\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}$ $\forall l=1,2,...,L$triangleright Gradient projection for layers
6:        end if
7:        $\mathbf{W}^{l}\leftarrow\mathbf{W}^{l}-\alpha\nabla_{\mathbf{W}^{l}}\mathcal{L}$triangleright Weight updates
8:     end while
9:     $\mathbf{X}^{l}_{t}\leftarrow\mathbf{X}^{l}_{t}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\mathbf{X}^{l}_{t}$triangleright Ensure uniqueness for new bases
10:     Perform SVD on $\mathbf{X}^{l}_{t}$ and put bases into $\mathcal{P}^{l}$triangleright Construct the first set of bases
11:     $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}\leftarrow\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}-\mathcal{P}^{l}\left(\mathcal{P}^{l}\right)^{T}\frac{\partial\mathbf{X}^{l}_{t}}{\partial\mathbf{X}_{t}}$
12:     Perform SVD on $\frac{\partial\mathbf{X}_{t}^{l}}{\partial\mathbf{X}_{t}}$ and put bases into $\mathcal{P}^{l}$triangleright Construct the second set of bases
13:  end for
@@ 4.1 Step
Baselines. For continual learning [11], in addition to SGD, which serves as a naive baseline using stochastic gradient descent to optimize the neural network, we adopt six algorithms cover three most important techniques in the field of continual learning: regularization – EWC [12] and SI [13], memory replay – GEM [14] and A-GEM [15], and gradient projection – OGD [1] and GPM [6]. The fundamental principle of each algorithm are outlined in Appendix B.3. For adversarial robustness, we adopt IGR [3] and AT [5].
We combine algorithms from fields of continual learning and adversarial robustness, such as EWC + IGR, to establish the baselines for robust continual learning. On the other hand, we apply the FGSM [5], PGD [16], and AutoAttack [17] to generate adversarial samples.
Figure 3: ACC varying with the number of learned tasks on datasets of Permuted MNIST (first row), Rotated MNIST (second row), CIFAR100 (third row) and miniImageNet (fourth row). ACC is measured on adversarial samples generated by AutoAttack (first column), PGD (second column) and FGSM (third column), as well as original samples (fourth column). The horizontal axis indicates the number of tasks learned by the neural network at present. The defense algorithm used here is IGR. Errors bars denote standard deviation.
@@
Benchmarks. We evaluate our approach on four supervised benchmarks. Permuted MNIST and Rotated MNIST are variants of MNIST dataset with 10 tasks applying random permutations of the input pixels and random rotations of the original images respectively [18, 19]. Split-CIFAR100 [13] is a random division of CIFAR100 into 10 subsets, each with 10 different classes. Split-miniImageNet is a random division of a part of the original ImageNet dataset [20] into 10 subsets, each with 5 different classes. All images for a specific class are exclusively present in one subset and no overlap of classes between subsets, thus these subsets can be considered as independent datasets, representing a sequence of 10 classification tasks.
@@
The values of $\frac{\partial\mathbf{x}^{2}}{\partial\mathbf{x}}$ (initial term in Eq.10) are solely determined by the weights of the first layer when feeding the same samples (see Fig.1). There are two options to make $\frac{\partial\mathbf{x}^{2}_{p,t}}{\partial\mathbf{x}_{p}}=\frac{\partial\mathbf{x}^{2}_{p,p}}{\partial\mathbf{x}_{p}}$: fixing the first layer after learning task $p$ or assigning an independent first layer to each task. The latter option is chosen for our experiments, as the former seriously diminishes the neural network’s learning ability in subsequent tasks. To ensure fairness, the same setup is applied to the baselines. Further details on architectures can be found in Appendix B.2.
Training details: For the MNIST experiments, the batch size, number of epochs, and input gradient regularization $\lambda$ are set to 32/10/50, respectively. For the Split-CIFAR100 experiments, the values are 10/100/1, and for the Split-miniImageNet experiments, they are 10/20/1. SGD is used as the optimizer. The hyperparameter configurations for adversarial attack and continuous learning algorithms are provided in Appendix B.3. All reported results are averaged over 5 runs with different seeds. We run the experiments on a local machine with three A800 GPUs.
Figure 4: As Fig. 3, but for defense algorithm Adversarial Training (AT) on PMNIST dataset. Here, we combine AT with continual learning methods GEM and GPM, which have shown superior ACC compared to other baselines in Fig. 3.
@@ 4.2.1 Adversarial Robustness
The results about robustness analysis on various datasets are presented in left three columns of Fig. 3, where different color lines represent the combinations by IGR with diverse continuous learning algorithms and DGP. Under attacks with increasing strengths (AutoAttack $>$ PGD $>$ FGSM), the proposed approach (orange lines) consistently exhibits a high level of effectiveness in maintaining robustness of neural networks enhanced by IGR. In contrast, the baseline such as IGR+GEM (purple lines), which performs well on MNIST datasets against PGD and FGSM attacks, demonstrates a significant decrease when fronted with AutoAttack. The advantage of our approach becomes even more evident when the number of learned tasks increases.
The results of maintaining the robustness enhanced by AT are presented in Fig. 4. The results further demonstrates that baselines fail to effectively maintain the robustness enhanced by AT against AutoAttack and PGD attacks after the neural network learns a sequence of new tasks, whereas GDP performs well. Compared to Fig. 3, the advantage of the proposed method than baselines is more pronounced in Fig. 4.
@@
Considering the collective insights presented in Figs. 3 and 4, it is crucial to underscore that the pursuit of an effective defense demands a tailored algorithm adept at accommodating variations in neural network’s parameters. Direct combinations of existing defense strategies and continual learning methods, as demonstrated in our experiments, fall short of achieving the desired goal of continuous robustness.
4.2.2 Classification Performance
@@
We also assess the ability of the proposed approach for continual learning (ACC on original samples), as illustrated in the fourth column of Fig. 3. Our DGP algorithm demonstrates comparable performance to GPM and GEM on datasets of Permuted MNIST and Rotated MNIST, effectively addressing the issue of catastrophic forgetting on these two datasets. However, results on the datasets of Split-CIFAR indicate that the performance of DGP is slightly inferior to GPM. We speculate that the reason for this could be that DGP stores a larger number of bases after each task than GPM, as DGP constrains the weight updates to be orthogonal to two sets of base vectors – one for stabilizing the final output (required in GPM) and another for stabilizing the sample gradients. Orthogonality to more base vectors restricts weight updates to a narrower subspace, thereby limiting the plasticity of the neural network. Overall, our approach effectively maintains adversarial robustness while exhibiting continual learning ability.
In addition, it is noteworthy that the performance curves of many well-known continual learning algorithms (e.g., EWC) closely approximate that of naive SGD (green lines). This important observation suggest a potential incompatibility between existing defense (here IGR) and continuous learning algorithms. The effectiveness of the latter can be significantly weakened when they are mixed into the training process. For instance, both EWC and IGR add a regularization term into the loss function, but their guidance on the direction of weight updates interferes with each other. Experimental results shown the incompatibility with another defense algorithm Distillation [21] are provided in Appendix B.4.
Figure 5: Gradient variation of samples from the first task $\mathcal{T}_{1}$ during continuous learning process trained with IGR. The variations are quantified through similarity.
@@ 4.2.3 Stabilization of sample gradients
Our approach maintains adversarial robustness by stabilizing the smoothness of sample gradients. To valid its stabilization effect, we record the variation of sample gradients on the first task during continuous learning process. Specifically, we randomly select $n$ samples at the end of learning $\mathcal{T}_{1}$ and compute their gradients related to correspondingly final outputs. After learning each new task $\mathcal{T}_{t}$, we recompute their gradients. The variation of gradients between $\mathcal{T}_{1}$ and $\mathcal{T}_{t}$ is quantified by similarity measure:
@@
where $\mathbf{g}$ is a flattened vector of sample gradients. The results on various datasets are presented in Fig. 5. The orange line (representing DGP) shows a relatively flat downward trend, demonstrating the proposed approach indeed has the effect of stabilizing the sample gradients of previous tasks, as the neural network’s weights update.
5 Related Works
One related work [22] also explores the emerging research direction of robust continual learning. A fundamental distinction between that work and ours is their approach requires partial data of previous tasks to be accessible, thereby focusing on the selection of a key subset of previous data and optimizing its utilization. In contrast, we follow the stricter yet realistic scenario in the field of continual learning that any data of the previous tasks cannot be revisited. Besides, the aim of our work is not to achieve stronger robustness on a single dataset, but rather to maintain robustness across multiple datasets encountered sequentially. For further insights into the advancements in adversarial robustness and continual learning, readers are encouraged to refer to dedicated surveys [2, 8].
@@ 6 Limitation and Discussion
In this work, we observe that the adversarial robustness gained by well-design defense algorithms is easily erased when the neural network learns new tasks. Direct combinations of existing defense and continuous learning algorithms fail to effectively address this issue, and may even give rise to conflicts between them. Therefore, we propose a novel gradient projection technique that can mitigate rapidly degradation of robustness in the face of drastic changes in model weights by collaborating with a class of defense algorithms through sample gradient smoothing.
According to our experiment, the proposed approach has certain limitations. First, as the number of base vectors becomes large, the stability of the neural network is enhanced to hold both robustness and performance across previous tasks. This stability may restrict the plasticity of the neural network, potentially reducing its ability to learn new tasks. Second, if there are numerous tasks and the matrix consisting of orthogonal bases reaches full rank, we approximate this matrix by performing SVD and selecting column vectors corresponding to a fraction of the largest singular values as the new orthogonal bases to free up rank space. Three, due to the extra challenge posed by our problem, the perturbation size of adversarial attacks under which the proposed method work effectively is slightly smaller than typical values in adversarial robustness literature (Please see Appendix B.3 for more details).